OpenStudy (anonymous):

(5x^2+2x+1)-(2x^2+x-3) how do ii find the difference ?

OpenStudy (anonymous):

Distribute he negative through: \[5x^2+2x+1-2x^2-x+3\] Group everything: \[(5x^2-2x^2)+(2x-x)+(1+3)=(5-3)x^2+(2-1)x+4=2x^2+x+4\]

OpenStudy (anonymous):

okayy thankszx

OpenStudy (anonymous):

No problem xD